how do i solve -4x-4y=4
-10x-7y=-2

Answer by stanbon(75887)

solve
-4x-4y=4
-10x-7y=-2
------------
Rearrange to get:
x-y = -1
---------------
x = y-1
------------
Substitute to solve for "y":
-10(y-1)-7y=-2
-10y+10-7y=-2
-17y = -12
y = 12/17
------------
Substitute into x=y-1 to solve for "x":
x = 12/17 - 1
x = -5/17
===============
Cheers,
Stan H.